Hydrolysis of C7H6Cl2 was carried out by refluxing with excess potassium hydroxide solution. The resulting solution was acidified with dilute nitric acid and excess silver nitrate added to precipitate the chloride ions as silver chloride,0.718g of silver chloride was formed from 0.805g of C7H6CL2. Calculate the number of chlorine atoms present in each molecule of C7H6CL2 which are liberated by hydrolysis. Suggest an explanation for the result obtained and give a possible structure of C7H6CL2


1
Expert's answer
2021-05-10T01:21:34-0400

"Mass = 0.805-0.718=0.087"


"Moles=\frac{0.087}{35.453}=2.45×10^{-3}"